Solve the following quadratic equations by completing the square.
x² - 6x -3 = 0

Answer by edjones(8007)

x^2-6x-3=0
x^2-6x =3 add 3 to both sides.
x^2-6x+9=3+9 take -6/2 and square it and add it to both sides making a perfect square on the left.
(x-3)^2=12 factor the left side.
x-3=+-sqrt(12) take sqrt of both sides.
x-3=+-sqrt(4*3)
x=3+-2sqrt(3)
Ed